OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

\[\frac{ 2 }{ 5(x-2) }=4x\] expand the denominator first \[\frac{ 2 }{ 5x - 10 } = 4x\] now u can bring the 5x-10 over to the RHS by reverse operation so we will have to times 4x. 2 = 4x(5x-10) 2 = 20x^2 - 40x 20x^2 - 40x - 2 = 0 divide everything by 2 10x^2 - 20x - 1 = 0

OpenStudy (anonymous):

u should be able to use a graphic calculator for this question so just use the calc to solve the last part.

OpenStudy (anonymous):

its 2 over 5 not 2 over 5 with the x and etc

OpenStudy (anonymous):